OpenStudy (anonymous):

Solve IVP t^2 dy/dt = 2cos^2 t , y(0)=pi/4

OpenStudy (anonymous):

\[t ^{2}\frac{ dy }{ dt }=2 \cos ^{2}t \] \[y(0)=\frac{ \Pi }{ 4 }\]

OpenStudy (anonymous):

no solution y(t) for all t>0 equals infinity since y(0)=pi/4 and y'(0)=infinity

OpenStudy (anonymous):

i'm confused sorry.

OpenStudy (anonymous):

gomen

OpenStudy (anonymous):

maybe the equation is supposed to be t^2 dy/dt = 2sin^2 t

OpenStudy (anonymous):

otherwise, at t=0, t^2 dy/dt = 2cos^2 t 0 * dy/dt = 2 dy/dt=2/0=infinity

OpenStudy (anonymous):

that only works if the initial condition is y(0)=-infinity

OpenStudy (anonymous):

i mainly need help with the integration part

OpenStudy (anonymous):

dy/dt=2cos^2 t/t^2 this can be expressed in terms of the sinintegral function Si(x)=integrate(sin(x)/x)

OpenStudy (anonymous):

one way is to reduce cos^2 t down to cos(2t) using the double angle identity, then use integration by parts to reduce the t^2 down to t

OpenStudy (anonymous):

that makes sense so then that gets rid of the 2 as well